A deck of cards has 3 blue, 4 black, and 6 purple cards. You pick 2 cards from the deck. Cards are not returned to the deck after they are picked. P(2 blue cards in a row) =
Answer by stanbon(75887)

A deck of cards has 3 blue, 4 black, and 6 purple cards. You pick 2 cards from the deck. Cards are not returned to the deck after they are picked. P(2 blue cards in a row) =
----
# of ways to pick 2 blue) = (3/13)(2/12) = 1/26
----------------------------------------------
Cheers,
Stan H.